# Enabling Approximate Joint Sampling in Diffusion LMs

### v1

@@ 3.3 Training Methodology
Unmasking Logic
In our discussion till now, we assumed the list of positions to be generated $\sigma$ to be given to us. For the current state-of-the-art diffusion LLMs, positions to be generated next are decided by an operator which is a function of the logits, denoted as TOP. Base diffusion language models specify their criterion to decide the most suitable position to generate next. For example, Dream models Ye et al. (2025) select the position with the least entropy to generate next, Llada Nie et al. (2025) unmasks position with the highest confidence. ADJUST is agnostic to the choice of the unmasking logic and uses the base model’s unmasking logic for generation. A subtle difference to note is that naive parallel sampling of $K$ tokens calls TOP on the same output embeddings $f({\bm{x}})$, $K$ times, yielding top $K$ positions with (say) least entropy in ${\bm{W}}f$. ADJUST uses the updated logits to get the index of the next position, $\sigma(k)=\text{TOP}\left\{{\bm{W}}{\bm{h}}_{k}\right\}$, which we qualitatively observe are often different than using the base model’s logits ${\bm{W}}f$.
Training Loss
We want our joint approximation $p_{\text{\text{ADJUST}}}$ to mimic the true joint distribution $p_{*}$. We do this by minimizing the KL divergence between the true joint and our approximation. Our training loss is hence a sum of KL divergence terms between $p^{\sigma(k)}\left(\cdot\mid{\bm{x}}\oplus\tilde{x}^{\sigma(1)}\oplus\ldots\oplus\tilde{x}^{\sigma(k-1)}\right)$ and $q^{\sigma(k)}\left(\cdot\mid{\bm{h}}_{k-1}\,,\,{\bm{x}}\oplus\tilde{x}^{\sigma(1)}\oplus\ldots\oplus\tilde{x}^{\sigma(k-1)}\right)$ for all values of $k\geq 2$. We still haven’t defined how $\bm{x}$} is generated or how we get $\sigma$ or $\bm{h}$} for computing the loss. We elaborate on these next.
Firstly, we describe how we select $\bm{x}$} for training. We first realize that $q$ would be evaluated at test on the distribution of partially masked strings encountered during denoising with the base diffusion model $p$. We consider $\bm{x}$} to be exactly the partially masked strings formed by the base model’s denoising process. That is, given an initial distribution of masked strings, we first sample an initial prompt string. We then run the base model’s denoising process to completion. Note that here we are using the base model’s unmasking logic and perform single token $K=1$ unmasking to sample from the true joint. The history of all partially masked strings encountered during the denoising process serve as our distribution of $\bm{x}$}. Following the same logic, we argue that $\sigma$ for a training sample $\bm{x}$} should be the next $K$ positions generated by the base model starting from $\bm{x}$} (on single token unmasking). Finally we describe how to get ${\bm{h}}_{k}$. Recall that, ${\bm{h}}_{k}$ is defined as $g\left({\bm{h}}_{k-1}\,,\,{\bm{x}}\oplus\tilde{x}^{\sigma(1)}\oplus\ldots\oplus\tilde{x}^{\sigma(k-1)}\right)$. We use the recursive formulation of $g$ during training as well to compute ${\bm{h}}_{k}$. Since $k$ is small ($\approx$ 4), this is not computationally expensive. The “unrolling” of $g$ during training is considered helpful by prior work  Liu et al. (2024a); Li et al. (2025) and hence we do not ablate on this choice. Putting these things together, we present our training algorithm in Alg 2.
@@
1:Input: Distribution $\mathcal{D}$ of input strings (each of which contains masked and unmasked positions), frozen discrete diffusion model $f,{\bm{W}}$ (where $\bm{W}$} is the language model head), number $K$ of tokens to generate in each step, base model’s unmasking logic TOP
@@
6:  ${\bm{h}}_{1},\ldots,{\bm{h}}_{K}\leftarrow f({\bm{x}}_{1}),\ldots,f({\bm{x}}_{K})$ $\triangleright$ Compute output embeddings
@@
11:  $\mathcal{L}(\theta)\leftarrow-\sum_{k}D_{\mathrm{KL}}({\bm{W}}\hat{{\bm{h}}}_{k},{\bm{W}}{\bm{h}}_{k})$
@@ 4.1 Experimental Setup
We consider two different state-of-the-art diffusion language models as base models for our empirical investigation. We consider a pretrained-only model (Dream-7B-Base) and an instruction-tuned model (Dream-7B-Instruct) Ye et al. (2025). Recall that, generating multiple tokens per denoising step leads to a deviation from the true underlying distribution. We therefore compare performances of different methods as a function of number of tokens $K$ generated for each denoising step. For simplicity of the setup, we generate the same number of tokens for each denoising step. We use the base model’s unmasking logic for our experiment (unless specified). Dream models’ unmasking logic chooses positions with the least entropy.
@@ Methods
We briefly describe the methods considered. 1) True joint sampling: This method corresponds to generating a single token $K=1$ in each denoising step. Recall that we established that generating one token per denoising step corresponds to sampling from the true joint distribution. Joint sampling serves as the oracle performance our method ADJUST aims to mimic. 2) Parallel sampling: This is the naive parallel sampling, default in diffusion models. Each of the $K$ generated tokens is sampled independently from the base model’s logits. 3) Energy-based model sampling Xu et al. (2024): EBM operates by drawing multiple token string unmasked completions (two in our experiments). It then chooses the completion with the lowest auto-regressive generative perplexity. The chosen string is then re-masked such that the final string has $K$ less masked positions (see Sec A.1 for discussion of hyper-parameter choices). We use the base model’s unmasking logic. Intuitively, the selection operator of EBM biases the samples towards the distribution of the auto-regressive model and away from the diffusion model. 4) Adaptive parallel decoding (APD)  Israel et al. (2025): is a specialized generation technique restricted to left-to-right decoding. At each denoising step, similar to EBM, APD samples a complete unmasked token string (in parallel). It then uses an auto-regressive evaluator (model) to compute the generative perplexity of all prefixes of the unmasked string and adaptively selects a prefix to commit and unmask. APD generates variable number of tokens in each denoising step. We compare with APD in Sec 4.4. 5) ADJUST (ours): This is our joint approximation algorithm (Sec 3) which updates the logits conditioned on the sampled tokens.
We acknowledge another baseline method discrete copula diffusion (DCD) Liu et al. (2024b) which we skip in our experimental setup. DCD uses an auto-regressive and diffusion model in tandem to get final output logits. Their formulation is orthogonal to our goal of mimic generation from the true joint distribution. (see Sec A.2 for additional discussion). We also highlight that all the baseline methods, assume an access to a “similarly” distributed auto-regressive model. This assumption though is valid for Dream models, can be restrictive. ADJUST does not rely on this assumption and works with any off-the-shelf diffusion model.
@@
We train ADJUST with a roll-out of three, that is, we assume $K=4$ for training. All models are trained for 2 epochs with constant learning rate of 5e-5 and batch size of 32. ADJUST has 200M parameters (compared to 7B parameters for base model). ADJUST assumes an initial distribution of prompts $\mathcal{D}$ (kindly refer Alg 2). For Dream-Base model we assume the initial prompt as just the beginning-of-sentence token, with max generation length of 128 tokens. We run the base model inference with a temperature of 1.0 for 100K samples to get our training dataset for Dream-Base. For Dream-Instruct, we subsample 20K unique questions from MetamathQA and generate with max generation length of 256 tokens, with temperature of 0.1 and nucleus of 0.9. We choose the data-generation parameters to align with our evaluation setup (though our results show robustness to these).
@@
Table 1: In this table we compare the NLL and the MAUVE scores for three different methods : Marginal, Joint and Energy-based (EBM). Across different number of tokens $K$ sampled per step, and different sampling temperature, joint sampling consistently outperforms both of the baselines, parallel sampling and EBM.
4.2 Unconditional Generation
We consider generation with the input prompt is just the beginning-of-sentence token, with a max generation length of 128. We call this task unconditional generation. We generate the output strings using our baseline methods and compare the negative log-likelihood (NLL) of the generated string as measured by an auto-regressive (AR) model (Qwen-2.5-7B for our setup). Recall that our target is to mimic the base model’s true joint distribution. Hence we also measure the generated strings MAUVE score Liu et al. (2021) (measured by GPT2-Large) with strings sampled from the true joint distribution, for given set of generation parameters. MAUVE score intuitively captures a notion of KL divergence between a two given sets of sentences and varies between zero and one with higher values implying a lower KL divergence between the distribution of the two sets.
Results
are presented in Table 1. We evaluate generations for two sampling temperatures, 0.6 and 1.0, highlighting the robustness of our method to sampling parameters. The MAUVE score is computed w.r.t. the samples from diffusion model’s true joint distribution, that is, samples generated with one token per diffusion step $K=1$ sampling. For $K=1$, EBM has lower NLL compared to our method, while (seemingly counterintuitively) having lower MAUVE. This is an expected behavior, as EBM samples two different completions for each denoising step and chooses the completion with a lower NLL. The selection of lower NLL completion biases the EBM model away from the true joint distribution of the diffusion model and more towards the underlying distribution of the auto-regressive model (used for computing NLL). This is also why EBM achieves lower NLL than parallel and ADJUST for $K=1$. We omit discussion of APD here and kindly refer readers to Sec 4.4.
4.3 Downstream Evaluation
We evaluate different methods on downstream tasks in this section. Sampling from the true joint distribution, $K=1$, leads to the best downstream accuracy. Hence here we show that our ADJUST method leads to better downstream accuracy than naive parallel sampling and closes matches performance of $K=1$ for any value of $K$. We use low temperature sampling (0.0-0.1) for generation. Exact sampling parameters are summarized in Sec D.1. We use lm-eval harness for evaluation. Due to computational constraints we limit our evaluation to 250 question from GSM8K & MBPP and 350 question (50 question per task) for Minerva math. Any mismatch between our numbers and the reported numbers (especially on Minerva-math) can be attributed to this artifact.
Results
Recall that the ADJUST for the base model is trained on domain-agnostic data (generations starting from begin-of-sentence token). Hence these results show the utility of a generically trained ADJUST for downstream tasks like GSM8K and MBPP in Table 2. Our method is better than marginal across all values of number of generated tokens in parallel even without ever specifically being prompted for math or code data. Since the downstream evaluation utilizes low temperature sampling, EBM leads to same performance as naive sampling. Hence we don’t report those numbers in the table. ADJUST for instruct model is trained on MetamathQA prompts. Hence we evaluate on math domain downstream tasks (GSM8K and Minerva-math). Across varying number of tokens sampled $K$ per denoising step, our joint sampling outperforms naive generation. We omit discussion of APD in this section and kindly refer readers to Sec 4.4.
K=1 K=2 K=3 K=4
GSM MBPP GSM MBPP GSM MBPP GSM MBPP
Base
Parallel 78.40 48.00 71.20 40.80 58.80 33.20 43.20 33.60
ADJUST(ours) 78.40 48.00 75.60 44.80 70.00 44.40 64.00 33.60
GSM MATH GSM MATH GSM MATH GSM MATH
Inst
Parallel 85.20 32.57 76.80 32.29 67.60 20.86 58.00 15.14
ADJUST(ours) 85.20 32.57 80.08 31.42 77.60 26.00 74.00 22.28
Runtime
We report the runtime numbers for naive sampling and our joint sampling in Fig 2. The left plot shows numbers for Dream-Base model on MAUVE metric, while the right plot shows numbers for Dream-Instruct on GSM8k. The numbers show tokens produced per second and show that our joint sampling method outperforms naive sampling for given throughput (tokens/sec) target. The discrepancy between tokens/sec between the two plots can be attributed to the longer sequence length and smaller batch-sizes used for GSM8k evaluation, as compared to unconditional generation. We report numbers for EBM on this figure, as EBM’s throughput (in our setup) is order of magnitude smaller than both joint and naive sampling. We report tokens/sec performance for unconditional generation using EBM in Table 3 ($\approx$ 50 toks/sec). We hypothesis that this is due to fact that EBM and APD both involve forward passes through multi layer (24 layers) models having thrice the number of parameters as our ADJUST model. Also note that ADJUST utilizes just a single transformer layer and hence is able to use parallel computing power of GPUs more effectively.
Figure 4: In this figure we compare adaptive parallel decoding (working left-to-right) to naive parallel decoding and ADJUST. Sampling $K>1$ tokens per diffusion step, for left-to-right decoding shows a steeper increase in NLL (left) than using the base model’s unmasking logic (see Table 4 for parallel generation, constrained to left-to-right decoding of $K>1$ tokens per step). Hence APD generation is not suited for such open-ended tasks and under-performs. The right figure shows GSM8k evaluation for high temperature sampling (temperature=1.0) for Dream-Instruct model. Left-to-right generation is naturally good for reasoning tasks, as shown by higher accuracy under APD when compared to naive/joint for $K=1$ tokens per step. APD leads to a quicker decrease in accuracy when compared to baseline and our method.
4.4 Comparison with adaptive parallel decoding
We first note that APD is a specialized left-to-right constrained decoding algorithm suited for reasoning heavy tasks evaluated at low temperatures. For e.g., for Dream-7B-Instruct, using APD we can generate (on an average) upto five tokens per denoising iteration for low temperature (=0.1) sampling on GSM8k prompts with negligible loss in performance (see Table 5). Here we investigate the performance of methods for open-ended tasks like unconditional generation and high temperature GSM8k sampling. We acknowledge the synthetic nature of high temperature GSM8k sampling and emphasize that these numbers are for completeness.
Results
They are encapsulated in caption of Fig 4. We highlight that we use left-to-right decoding for APD and use base model’s unmasking logic for parallel and ADJUST. The results show that ADJUST is applicable across tasks and sampling temperatures where APD fails.
Appendix A Additional Discussions
A.1 Hyper-parameter selection
EBM
We use Qwen-2.5-0.5B to evaluate generative perplexity of sampled strings and sample two string completions. For EBM we choose just two samples as this was shown to be optimal number presented in the EBM paper Xu et al. (2024). We ablated with 4 choices as well but got similar results. Hence we just consider sampling 2 options. Another consideration for EBM is the choice of the auto-regressive model. We take the model to be Qwen2.5-0.5B. We also tried Qwen2.5-7B but that also didn’t help improve the performance much. The numbers are reported as EBM (Big) in Table 3.
A.2 Discrete Copula Diffusion
DCD uses an auto-regressive model of the same size and distribution as the base diffusion model to generate a complete unmasked token string. An auto-regressive generation strategy uses $|\mathcal{M}({\bm{x}})|$ forward passes of the AR model (of the same size as the diffusion model). We can alternatively use $|\mathcal{M}({\bm{x}})|$ forward passes of the diffusion model (with $K=1$) to directly sample from the true joint (without using the AR model).
DCD Liu et al. (2024b) aims to construct a joint distribution over the whole token string instead of just tokens to be unmasked. Since they have a joint over all the masked positions, their method works in a regime of aggressive unmasking. ADJUST does auto-regressive generation over only the tokens to be unmasked. Hence, our method (intuitively) predicts a joint over positions to be unmasked and not the whole sentence. This makes our method much more computationally friendly then DCD.
Appendix B Additional Experiments
You may include other additional sections here.
B.1 EBM Experiments

### v2

@@ 3.3 Training Methodology
We want our joint approximation $p_{\text{\text{ADJUST}}}$ to mimic the true joint distribution $p_{*}$. We do this by minimizing the KL divergence between the true joint and our approximation. Our training loss is hence a sum of KL divergence terms between $p\left(\cdot\mid{\bm{x}}\oplus\tilde{x}^{\sigma(1)}\oplus\ldots\oplus\tilde{x}^{\sigma(k-1)}\right)$ and $q\left(\cdot\mid{\bm{h}}_{k-2}\,,\,{\bm{x}}\oplus\tilde{x}^{\sigma(1)}\oplus\ldots\oplus\tilde{x}^{\sigma(k-1)}\right)$ for all values of $k\geq 2$. Note that the probabilities represent logits for each position in the sequence. We still haven’t defined how $\bm{x}$} is generated or how we get $\sigma$ or $\bm{h}$} for computing the loss. We elaborate on these next.
Firstly, we describe how we select $\bm{x}$} for training. We first realize that $q$ would be evaluated at test on the distribution of partially masked strings encountered during denoising with the base diffusion model $p$. We consider $\bm{x}$} to be exactly the partially masked strings formed by the base model’s denoising process. That is, given an initial distribution of masked strings, we first sample an initial prompt string. We then run the base model’s denoising process to completion. Note that here we are using the base model’s unmasking logic (see App E for details) and perform single token $K=1$ unmasking to sample from the true joint. The history of all partially masked strings encountered during the denoising process serve as our distribution of $\bm{x}$}. Following the same logic, we argue that $\sigma$ for a training sample $\bm{x}$} should be the next $K$ positions generated by the base model starting from $\bm{x}$} (on single token unmasking). Finally we describe how to get ${\bm{h}}_{k}$. Recall that, ${\bm{h}}_{k}$ is defined as $g\left({\bm{h}}_{k-1}\,,\,{\bm{x}}\oplus\tilde{x}^{\sigma(1)}\oplus\ldots\oplus\tilde{x}^{\sigma(k)}\right)$. We use the recursive computation of embeddings with teacher forcing of token inputs during training. The “unrolling” of $g$ during training is considered helpful by prior work  (Liu et al., 2024a; Li et al., 2025) and hence we do not ablate on this choice. Putting these things together, we present our training algorithm in Alg 2.
@@
1:Input: Distribution $\mathcal{D}$ of input strings (each of which contains masked and unmasked positions), frozen discrete diffusion model $f,{\bm{W}}$ (where $\bm{W}$} is the language model head), number $K$ of tokens to generate in each step
@@
6:  ${\bm{h}}_{1},\ldots,{\bm{h}}_{K}\leftarrow f({\bm{x}}_{1}),\ldots,f({\bm{x}}_{K})$
@@
11:  $\mathcal{L}(\theta)\leftarrow$ KL divergence between ${\bm{W}}{\bm{h}}_{k}$ and ${\bm{W}}\hat{{\bm{h}}}_{k}$
@@ 4 Experiments
K=1 K=2 K=4 K=8
NLL Mauve NLL Mauve NLL Mauve NLL Mauve
T=0.6 Parallel 1.11 1.00 1.61 0.89 2.21 0.29 2.21 0.02
EDLM 0.83 0.98 1.33 0.94 1.89 0.31 2.05 0.03
ADJUST (ours) 1.11 1.00 1.23 0.97 1.64 0.79 1.66 0.67
T=1.0 Parallel 4.25 1.00 5.68 0.76 6.52 0.31 7.10 0.19
EDLM 3.75 0.96 5.26 0.83 6.18 0.39 6.56 0.16
ADJUST (ours) 4.25 1.00 4.99 0.96 5.58 0.87 5.91 0.84
Table 1: In this table we compare the NLL and the MAUVE scores for three different methods : Marginal, Joint and Energy-based (EDLM). The base diffusion model is Dream-Base-7B. Across different number of tokens $K$ sampled per step, and different sampling temperature, joint sampling consistently outperforms both of the baselines, parallel sampling and EDLM.
@@ 4.1 Experimental Setup
We consider four different state-of-the-art discrete diffusion language models as base models for our empirical investigation. We consider a pretrained-only model (Dream-7B-Base, Llada-7B-Base) and an instruction-tuned model (Dream-7B-Instruct, Dream-7B-Coder) (Ye et al., 2025). While ADJUST can be use to obtain more tighter ELBO numbers, we focus our evaluation on solely the generative performance of different methods. Recall that, generating multiple tokens per denoising step leads to a deviation from the true underlying distribution. We therefore compare performances of different methods as a function of number of tokens $K$ generated for each denoising step. For simplicity of the setup, we generate the same number of tokens for each denoising step. We use the base model’s unmasking logic for our experiment (unless specified). Dream models’ unmasking logic chooses positions with the least entropy and Llada uses the most confident token.
@@ Methods
We briefly describe the methods considered. 1) True joint sampling: This method corresponds to generating a single token $K=1$ in each denoising step. Recall that we established that generating one token per denoising step corresponds to sampling from the true joint distribution. Joint sampling serves as the oracle performance our method ADJUST aims to mimic. 2) Parallel sampling: This is the naive parallel sampling, default in diffusion models. Each of the $K$ generated tokens is sampled independently from the base model’s logits. 3) Energy-based model sampling (Xu et al., 2024): EDLM operates by drawing multiple token string unmasked completions (two in our experiments). It then chooses the completion with the lowest auto-regressive generative perplexity. The chosen string is then re-masked such that the final string has $K$ less masked positions (see Sec A.1 for discussion of hyper-parameter choices). We use the base model’s unmasking logic. 4) ADJUST (ours): This is our joint approximation algorithm (Sec 3) which updates the logits conditioned on the sampled tokens.
We acknowledge other baseline methods like adaptive parallel decoding(APD) (Israel et al., 2025) and discrete copula diffusion (DCD) (Liu et al., 2024b) which we omit from the main text. Their formulation is orthogonal to our goal of mimic generation from the true joint distribution. (see Sec B for results and discussion). We also highlight that all the baseline methods, assume an access to a “similarly” distributed auto-regressive model. This assumption though is valid for Dream models, can be restrictive. ADJUST does not rely on this assumption and works with any off-the-shelf diffusion model.
@@
We train ADJUST with a roll-out of three, that is, we assume $K=4$ for training. All models are trained for 2 epochs with constant learning rate of 5e-5 and batch size of 32. ADJUST has 200M parameters (compared to 7B parameters for base model). Recall that ADJUST assumes an initial distribution of prompts $\mathcal{D}$ (kindly refer Alg 2). For pretrained-only models we assume the initial prompt as just the beginning-of-sentence token, with max generation length of 128 tokens. We run the base model inference with a temperature of 1.0 for 100K samples to get our training dataset. For Dream-Instruct, we subsample 20K unique questions from MetamathQA (Yu et al., 2023) and generate with max generation length of 256 tokens, with temperature of 0.1 and nucleus of 0.9. For Dream-Coder we similarly subsample 20K prompts from Opencoder-LLM’s opc-sft-stage2 (Huang et al., 2024). All data-generation runs take one GPU day on H100. All training runs take four GPU days on H100. Note that we are only sampling the prompts and not the ground truth responses.
Model Task K=1 K=2 K=4 K=8
Parallel ADJ. Parallel ADJ. Parallel ADJ.
Llada-Base GSM 68.94 60.61 66.67 54.55 57.58 19.70 28.79
MBPP 40.40 28.40 37.40 19.40 29.40 6.60 20.00
HEval 33.54 25.00 29.88 12.80 23.78 5.49 10.37
Dream-Base GSM 78.40 71.20 75.60 43.20 64.00 11.60 36.80
MBPP 48.00 40.80 44.80 33.60 33.60 18.40 19.60
Dream-Inst GSM 85.20 76.80 80.08 58.00 74.00 20.00 44.40
MATH 32.57 32.29 31.42 15.14 22.28 1.14 5.70
Dream-Coder MBPP 63.00 47.00 57.40 12.60 16.00 4.60 9.20
HEval 75.00 55.49 73.17 20.73 57.32 1.83 28.05
4.2 Unconditional Generation
We consider generation with the input prompt as the beginning-of-sentence token, with a max generation length of 128. This unconditional generation task aims to measure the language modeling capability of the model. We generate the output strings using our baseline methods and compare the negative log-likelihood (NLL) of the generated string as measured by an auto-regressive (AR) model (Qwen-2.5-7B for our setup). Recall that our target is to mimic the base model’s true joint distribution. Hence we also measure the generated strings MAUVE score (Liu et al., 2021) (measured by GPT2-Large) with strings sampled from the true joint distribution, for given set of generation parameters. MAUVE score intuitively captures a notion of KL divergence between a two given sets of sentences and varies between zero and one with higher values implying a lower KL divergence between the distribution of the two sets.
Results
are presented in Table 1. We present results for pre-trained only Dream-Base-7B. We evaluate generations for two sampling temperatures, 0.6 and 1.0, highlighting the robustness of our method to sampling parameters. The MAUVE score is computed w.r.t. the samples from diffusion model’s true joint distribution, that is, samples generated with one token per diffusion step $K=1$ sampling. For $K=1$, EDLM has lower NLL compared to our method, while (seemingly counterintuitively) having lower MAUVE. This is an expected behavior, as EDLM samples two different completions for each denoising step and chooses the completion with a lower NLL. The selection of lower NLL completion biases the EDLM model away from the true joint distribution of the diffusion model and more towards the underlying distribution of the auto-regressive model (used for computing NLL). This is also why EDLM achieves lower NLL than parallel and ADJUST for $K=1$. Results for $K=8$ show how well ADJUST generalizes beyond the training rollouts.
4.3 Downstream Evaluation
Here we show that our ADJUST method leads to better downstream accuracy than naive parallel sampling and closes matches performance of $K=1$ for any degree of parallel generation. We use low temperature sampling (0.0-0.1) for generation. Due to computational constraints we limit our evaluation to 250 question from GSM8K & MBPP and 350 question (50 question per task) for Minerva math. Exact sampling and evaluation parameters are summarized in Sec D.1.
Results
Recall that the ADJUST for the pretrained-only model is trained on domain-agnostic data (generations starting from begin-of-sentence token). Hence the results for Dream-Base and Llada-Base show the utility of a generically trained ADJUST for downstream tasks like GSM8K and MBPP in Table 2. Our method is better than marginal across all values of number of generated tokens in parallel even without ever specifically being prompted for math or code data. Since the downstream evaluation utilizes low temperature sampling, EDLM leads to same performance as naive sampling. Hence we don’t report those numbers in the table. ADJUST for Dream-Instruct model is trained on MetamathQA prompts. Hence we evaluate on math domain downstream tasks (GSM8K and Minerva-math). Similarly, since we train Dream-Coder on opc-sft-stage2 prompts, we evaluate it on coding tasks. Results for $K=8$ lie outside the training rollout range of $K=4$ and demonstrate the generalization of ADJUST along longer prediction horizons.
Runtime
We report the runtime numbers for naive sampling and our joint sampling in Fig 2. The left plot shows numbers for Dream-Base model on MAUVE metric, while the right plot shows numbers for Dream-Instruct on GSM8k. The numbers show tokens produced per second and show that our joint sampling method outperforms naive sampling for given throughput (tokens/sec) target. The discrepancy between tokens/sec between the two plots can be attributed to the longer sequence length and smaller batch-sizes used for GSM8k evaluation, as compared to unconditional generation. We report tokens/sec performance for unconditional generation using EDLM in Table 6 ($\approx$ 50 toks/sec). We hypothesize that this is due to fact that EDLM and APD both involve forward passes through multi layer (24 layers) models, while ADJUST utilizes just a single transformer layer and is thus much faster.
4.4 ParallelBench
ParallelBench (Kang et al., 2025) is a recent benchmark that quantifies the quality degradation under parallel decoding with diffusion LMs. We here report numbers for a representative task word-to-sentence using Dream-Base-7B as the base diffusion model in Table 3. The task prompt asks the model to generate a coherent sentence with a set of given words present in the sentence. A generation is valid if it’s both coherent (grammatically) and includes all the specified words. We report the fraction of valid generation for three task difficulty levels (easy (E), medium (M) and hard (H)). We kindly refer the reader to App. Table 5) for results on the other tasks in parallelbench.
W2S(E) W2S(M) W2S(H)
True Joint 82.0 87.0 69.0
Parallel (K=4) 81.0 77.0 63.0
ADJUST (K=4) 84.0 85.0 69.0
Table 3: In this table, we report numbers on the word-to-sentence (W2S) task of ParallelBench (Kang et al., 2025).
5 Conclusion
We study the problem of sampling multiple tokens for each denoising step of discrete diffusion language models. We first showed that naively sampling more than one from the model’s output logits deviates from the underlying distribution of the diffusion model. Our main contribution was the formalization of our approximate joint sampler ADJUST, along with proposing a new training methodology based on sampling trajectory. For future work, we plan to explore the whether ADJUST can speed up pre-training diffusion backbone or lead to a better pre-trained diffusion model, following prior work on multi-token prediciton for auto-regressive models.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Additional Discussions
A.1 Hyper-parameter selection
EDLM
We use Qwen-2.5-0.5B to evaluate generative perplexity of sampled strings and sample two string completions. For EDLM we choose just two samples as this was shown to be optimal number presented in the EDLM paper (Xu et al., 2024). We ablated with 4 choices as well but got similar results. Hence we just consider sampling 2 options. Another consideration for EDLM is the choice of the auto-regressive model. We take the model to be Qwen2.5-0.5B. We also tried Qwen2.5-7B but that also didn’t help improve the performance much. The numbers are reported as EDLM (Big) in Table 6.
Appendix B Additional Baselines
B.1 Discrete Copula Diffusion
DCD uses an auto-regressive model of the same size and distribution as the base diffusion model to generate a complete unmasked token string. An auto-regressive generation strategy uses $|\mathcal{M}({\bm{x}})|$ forward passes of the AR model (of the same size as the diffusion model). We can alternatively use $|\mathcal{M}({\bm{x}})|$ forward passes of the diffusion model (with $K=1$) to directly sample from the true joint (without using the AR model). DCD Liu et al. (2024b) aims to construct a joint distribution over the whole token string instead of just tokens to be unmasked. Since they have a joint over all the masked positions, their method works in a regime of aggressive unmasking. ADJUST does auto-regressive generation over only the tokens to be unmasked. Hence, our method (intuitively) predicts a joint over positions to be unmasked and not the whole sentence. This makes our method much more computationally friendly then DCD.
B.2 Adaptive parallel decoding
(APD)  (Israel et al., 2025): is a specialized generation technique restricted to left-to-right decoding. At each denoising step APD samples a complete unmasked token string (in parallel). It then uses an auto-regressive evaluator (model) to adaptively selects a prefix string to commit and unmask. Because APD generates variable number of tokens in each denoising step and is strictly left-to-right.
We first note that APD is a specialized left-to-right constrained decoding algorithm suited for reasoning heavy tasks evaluated at low temperatures. For e.g., for Dream-7B-Instruct, using APD we can generate (on an average) upto five tokens per denoising iteration for low temperature (=0.1) sampling on GSM8k prompts with negligible loss in performance. Here we investigate the performance of methods for open-ended tasks like unconditional generation and high temperature GSM8k sampling. We acknowledge the synthetic nature of high temperature GSM8k sampling and emphasize that these numbers are for completeness.
Results are encapsulated in caption of Fig 4. We highlight that we use left-to-right decoding for APD and use base model’s unmasking logic for parallel and ADJUST. The results show that ADJUST is applicable across tasks and sampling temperatures where APD fails.
Figure 4: In this figure we compare adaptive parallel decoding (working left-to-right) to naive parallel decoding and ADJUST. Sampling $K>1$ tokens per diffusion step, for left-to-right decoding shows a steeper increase in NLL (left) than using the base model’s unmasking logic. Hence APD generation is not suited for such open-ended tasks and under-performs. The right figure shows GSM8k evaluation for high temperature sampling (temperature=1.0) for Dream-Instruct model. Left-to-right generation is naturally good for reasoning tasks, as shown by higher accuracy under APD when compared to naive/joint for $K=1$ tokens per step. APD leads to a quicker decrease in accuracy when compared to baseline and our method.
Appendix C Additional Experiments from omitted from the main text
C.1 Complete NLL and MAUVE numbers
See Table 4 for all numbers
@@ C.1 Complete NLL and MAUVE numbers
K=5 K=6 K=7 K=8
NLL Mauve NLL Mauve NLL Mauve NLL Mauve
T=0.6 Parallel 2.00 0.12 2.11 0.07 2.19 0.05 2.21 0.02
EDLM 1.87 0.14 1.95 0.07 2.03 0.05 2.05 0.03
ADJUST (ours) 1.47 0.79 1.54 0.77 1.60 0.76 1.66 0.67
T=1.0 Parallel 6.78 0.22 6.80 0.24 6.89 0.17 7.10 0.19
EDLM 6.19 0.27 6.26 0.26 6.35 0.21 6.56 0.16
ADJUST (ours) 5.67 0.87 5.76 0.85 5.83 0.85 5.91 0.84
@@
Table 4: In this table we compare the NLL and the MAUVE scores for three different methods : Marginal, Joint and Energy-based (EDLM). Across different number of tokens $K$ sampled per step, and different sampling temperature, joint sampling consistently outperforms both of the baselines, parallel sampling and EDLM.
C.2 Complete ParallelBench Numbbers
See Table 5
Model W2S(E) W2S(M) W2S(H) Copy Ins(idx) Ins(rand) Rm(idx)
Dream (x1) 82.0 87.0 69.0 73.0 30.0 42.0 72.0
Parallel (x4) 81.0 77.0 63.0 55.0 12.0 13.0 57.0
ADJUST (x4) 84.0 85.0 69.0 63.0 21.0 20.0 62.0
Model Rm(rand) Rpl(idx) Rpl(rand) Reverse Shuffle Sort
Dream (x1) 62.0 51.0 52.0 68.0 62.0 21.0
Parallel (x4) 36.0 35.0 1.0 27.0 24.0 13.0
ADJUST (x4) 39.0 46.0 3.0 39.0 14.0 16.0
Table 5: Evaluation on ParallelBench using Dream-Base-7B model
C.3 EDLM Experiments